OpenStudy (anonymous):

Please help If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22

OpenStudy (anonymous):

The first step is to rewrite the first equation in terms of y You can do this by adding y to both sides and adding 9 to both sides: 8x - y = -9 => y = 8x + 9 Now plug this value for y into the second equation since you know that 4x - 3y = -22 and that y = 8x + 9 you can say that 4x - 3(8x + 9) = -22 now you just have to simplify and solve